# Higher-Order Reverse Isoperimetric Inequalities for Log-concave Functions

### v4

@@ 1. Introduction
$\quad\quad\quad\quad\left.\text{upper-semi-continuous, and integrable}.\right\}$
@@
We write $\mathrm{BV}(\mathbb{R}^{n})$ to denote the space of functions of bounded variation on $\mathbb{R}^{n}$. We mention here that given a function in $\mathrm{BV}(\mathbb{R}^{n})$, there exists a vector-valued measure on $\mathbb{R}^{n}$ associated to it, its so-called variation measure. One has that $\operatorname{LC}_{n}\subset\mathrm{BV}(\mathbb{R}^{n})$. Recently, Rotem [63, Proposition 2.5] established a formula for the variation measure of $f\in\operatorname{LC}_{n}$; for a self-contained study of log-concave functions, we choose to use this explicit representation in our definitions. We recall that, for $\in\operatorname{LC}_{n}, f\nabla f$ exists almost everywhere on $\mathbb{R}^{n}$. Moreover, $\mathop{\mathrm{supp}}\nolimits(f)$ is convex and, therefore, $\partial\mathop{\mathrm{supp}}\nolimits(f)$ admits $\mathcal{H}^{n-1}$-almost everywhere an outer-unit normal.
@@
We address the existence and uniqueness of the asymmetric LYZ body in Section 2.1.
@@
Concerning the asymptotics for $p\to\infty$, we will see that $R_{\infty}^{m}f$ may be a compact, convex set, but it may also be $\mathbb{R}^{nm}$. Additionally, the limit $\lim_{p\to\infty}\Gamma\left(1+p\right)^{-\frac{1}{p}}R^{m}_{p}f$ will always be a compact, convex set, but it may be a singleton containing the origin, as the following statement shows.
@@
Let $K\subset\mathbb{R}^{n}$ be a convex body and set $\gamma_{K}(x)=e^{-\frac{\|x\|_{K}^{2}}{2}}$. Then, for $p>0$
$R^{m}_{p}\gamma_{K}=\sqrt{2}\left(2^{\frac{n(m-1)}{2}}\frac{\Gamma\left(1+\frac{nm+p}{2}\right)}{\Gamma\left(1+\frac{n}{2}\right)}\right)^{\frac{1}{p}}R_{p}^{m}K.$
@@
Let $f\in\operatorname{LC}_{n}$. Then, for $1<p<q<\infty$,
@@ 1.4. Functional, higher-order Rogers-Shephard inequality
Let $f\in\operatorname{LC}_{n}$. Then, $\mathop{\mathrm{supp}}\nolimits(g_{f,m})\subseteq\mathbb{R}^{nm}$ contains the origin in its interior and
@@ 2.1. The LYZ body of log-concave functions
In usual applications in the literature (compare [50, 68, 49, 18]), the LYZ body of $f$ describes the unique origin symmetric convex body whose surface area measure equals the even part of $\mu_{f}$. The symmetric LYZ body is a fundamental tool in establishing affine-Sobolev-type inequalities [73, 52, 36, 68, 51, 34]. Similar applications of the asymmetric body $\langle f\rangle$ can be found in [32, 33]. We emphasize that, when dealing with $f\in\operatorname{LC}_{n}$, the body $\langle f\rangle$ is not necessarily symmetric, since the level sets of $f$ are always convex and bounded and, therefore, the measure $\mu_{f}$ we retrieve in Proposition 2.3 is centered by construction, without requiring any further step.
@@ 3. Sets associated with functions
This section is dedicated to obtaining (1.25), (1.27), and Theorem 1.6 as special cases of a more general statement. We begin by recalling that a function $f:\mathbb{R}^{n}\longrightarrow\mathbb{R}_{+}$ is $s$-concave, $s>0$, if it is upper-semi continuous, not identically zero, and, for $\lambda\in[0,1]$ and $x,y\in\mathbb{R}^{n}$ such that $f(x)f(y)>0$, one has
@@
(3.13) $\begin{split}\rho_{K_{p}(g)}(\theta)&=\left(p\mathcal{M}_{\widetilde{g(\cdot\theta)}}(p)\right)^{\frac{1}{p}}\\ &=\begin{cases}\left(\frac{p}{\|g\|_{\infty}}\int_{0}^{\infty}g(r\theta)r^{p-1}\,\mathrm{d}r\right)^{\frac{1}{p}},&p>0,\\ \exp\left(\frac{1}{\|g\|_{\infty}}\int_{0}^{\infty}(-\frac{\partial}{\partial r}g(r\theta))\log(r)\;\mathrm{d}r\right),&p=0,\\ \left(\frac{p}{\|g\|_{\infty}}\int_{0}^{\infty}r^{p-1}(g(r\theta)-g(0))\;\mathrm{d}r\right)^{\frac{1}{p}},&p\in(-1,0).\end{cases}\end{split}$
@@
When $p\in(-1,0)$, an analogous requirement to (3.14) is that $g$ has maximum at the origin and the function $x\mapsto g(0)-g\left(\frac{x}{|x|^{2}}\right)$ has finite (p)(-p)-moment. Indeed, from a change of variables
$p\int_{0}^{\infty}r^{p-1}(g(r\theta)-g(0))\mathrm{d}\,r=|p|\int_{0}^{\infty}r^{|p|-1}\left(g(0)-g\left(\frac{\theta}{r}\right)\right)\mathrm{d}\,r,$
@@
$K_{\infty}(g):=\lim_{p\to+\infty}K_{p}(g)=\mathop{\mathrm{supp}}\nolimits(g),$
@@ 4.2. The Functional Covariogram
Let $f\in\operatorname{LC}_{n}$. Then, for every $\bar{\theta}\in\mathbb{S}^{nm-1}$,
@@
Let $f\in\operatorname{LC}_{n}$. Then, $g_{f,m}$ is log-concave and, for a fixed $\bar{x}\in\mathbb{R}^{nm}$,
@@ 4.3. The approach by radial mean bodies
Clearly, the definition is well-posed, since the restriction of a summable log-concave function on $\mathbb{R}^{n}$ to a line is a summable one-dimensional log-concave function. Define the function $f_{y,m}:\mathbb{R}^{nm}\longrightarrow\mathbb{R}_{+}$, by
$f_{y,m}(\bar{x})=\min_{0\leq i\leq}\left\{f(y-x_{i})\right\},$
@@
We finish this section by characterizing the radial mean bodies of a particular class of functions which contains those of Corollary 1.5.
@@
Let $\varphi:\mathbb{R}_{+}\longrightarrow\mathbb{R}_{+}$ be a differentiable function with maximum at 0 that is decreasing to zero at $\infty$. Fix $K\subset\mathbb{R}^{n}$ a convex body and define the function $f:\mathbb{R}^{n}\longrightarrow\mathbb{R}_{+}$ by $f=\varphi(\|\cdot\|_{K})$. Then, for $p>0$,
$R^{m}_{p}f=\left(\frac{\int_{0}^{\infty}(-\varphi^{\prime}(t))t^{nm+p}\mathrm{d}t}{n\int_{0}^{\infty}\varphi(t)t^{n-1}\;\mathrm{d}t}\right)^{\frac{1}{p}}R_{p}^{m}K.$
@@
First, we compute $g_{f,m}$. For every $\bar{x}\in\mathbb{R}^{nm},$ it holds
$g_{f,m}(\bar{x})$ $=\int_{\mathbb{R}^{n}}\min_{0\leq i\leq m}\varphi(\|y-x_{i}\|_{K})\;\mathrm{d}y=\int_{\mathbb{R}^{n}}\min_{0\leq i\leq m}\int_{\|y-x_{i}\|_{K}}^{\infty}(-\varphi^{\prime}(t))\;\mathrm{d}t\;\mathrm{d}y$
$=\int_{0}^{\infty}(-\varphi^{\prime}(t))\mathrm{vol}_{nm}\left(tK\cap\bigcap_{i=1}^{m}\left(tK+x_{i}\right)\right)\;\mathrm{d}t$
$=\int_{0}^{\infty}(-\varphi^{\prime}(t))t^{nm}\mathrm{vol}_{nm}\left(K\cap\bigcap_{i=1}^{m}\left(K+t^{-1}x_{i}\right)\right)\;\mathrm{d}t$
$=\int_{0}^{\infty}(-\varphi^{\prime}(t))t^{nm}g_{K,m}(t^{-1}\bar{x})\;\mathrm{d}t.$
@@
$\int_{\mathbb{R}^{n}}\varphi(\|x\|_{K})\;\mathrm{d}x$ $=\int_{\mathbb{S}^{n-1}}\int_{0}^{\infty}\varphi(r\|\theta\|_{K})r^{n-1}\,\mathrm{d}r\;\mathrm{d}\theta$
$=\left(n\int_{0}^{\infty}\varphi(r)r^{n-1}\;\mathrm{d}r\right)\left(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\|\theta\|_{K}^{-n}\;\mathrm{d}\theta\right)$
$=\mathrm{vol}_{n}(K)\left(n\int_{0}^{\infty}\varphi(r)r^{n-1}\;\mathrm{d}r\right).$
We compute the Dirichlet-type integral involving $g_{f,m}$ and obtain, for every $\bar{\theta}\in\mathbb{S}^{nm-1}$,
@@
$=\frac{p}{\|f\|_{1}}\int_{0}^{\infty}\int_{0}^{\infty}(-\varphi^{\prime}(t))t^{nm}g_{K,m}(rt^{-1}\bar{\theta})r^{p-1}\;\mathrm{d}t\;\mathrm{d}r$
$=\frac{p}{\|f\|_{1}}\int_{0}^{\infty}\int_{0}^{\infty}(-\varphi^{\prime}(t))t^{nm+p}g_{K,m}(r\bar{\theta})r^{p-1}\;\mathrm{d}t\;\mathrm{d}r$
$=\left(\frac{p}{\mathrm{vol}_{n}(K)}\int_{0}^{\infty}g_{K,m}(r\bar{\theta})r^{p-1}\mathrm{d}r\right)\left(\frac{\mathrm{vol}_{n}(K)}{\|f\|_{1}}\int_{0}^{\infty}(-\varphi^{\prime}(t))t^{nm+p}\mathrm{d}t\right)$
$=\rho_{R^{m}_{p}K}(\bar{\theta})^{p}\left(\frac{\int_{0}^{\infty}(-\varphi^{\prime}(t))t^{nm+p}\mathrm{d}t}{n\int_{0}^{\infty}\varphi(t)t^{n-1}\;\mathrm{d}r}\right).$
The claim follows. ∎
We remark that, if $f=e^{-\|\cdot\|_{K}},$ i.e. $\varphi(t)=e^{-t}$, we obtain
$R^{m}_{p}f=\left(\frac{\Gamma(nm+p+1)}{\Gamma(n+1)}\right)^{\frac{1}{p}}R_{p}^{m}K.$
Using $\Gamma(n+1)^{-\frac{1}{p}}\to 1$ as $p\to\infty$, we have
@@
$=\lim_{p\to\infty}\left(\frac{\Gamma(nm+p+1)}{\Gamma(p+1)}\right)^{\frac{1}{p}}R_{p}^{m}K=R_{\infty}^{m}K=D^{m}(K),$
since, by Stirling’s approximation,
$\left(\frac{\Gamma(nm+p+1)}{\Gamma(p+1)}\right)^{\frac{1}{p}}\xrightarrow{p\to\infty}1.$
Corollary 1.5 then follows from Theorem 4.7 with the choice $\varphi(t)=e^{-\frac{t^{2}}{2}}$ and an application of Stirling’s approximation.

### v5

@@ 1. Introduction
$\quad\quad\quad\quad\left.\text{upper semi-continuous, and integrable}.\right\}$
@@
We write $\mathrm{BV}(\mathbb{R}^{n})$ to denote the space of functions of bounded variation on $\mathbb{R}^{n}$. We mention here that given a function in $\mathrm{BV}(\mathbb{R}^{n})$, there exists a vector-valued measure on $\mathbb{R}^{n}$ associated to it, its so-called variation measure. One has that $\operatorname{LC}_{n}\subset\mathrm{BV}(\mathbb{R}^{n})$. Recently, Rotem [63, Proposition 2.5] established a formula for the variation measure of $f\in\operatorname{LC}_{n}$; for a self-contained study of log-concave functions, we choose to use this explicit representation in our definitions. We recall that, for $\in\operatorname{LC}_{n}, f\nabla f$ exists almost everywhere on $\mathbb{R}^{n}$. Moreover, $\mathop{\mathrm{supp}}\nolimits(f)$ is convex and, therefore, $\partial\mathop{\mathrm{supp}}\nolimits(f)$ admits $\mathcal{H}^{n-1}$-almost everywhere an outer-unit normal. A fundamental instrument in our investigation is the LYZ body, originally introduced by Lutwak, Yang, and Zhang in [52]. We present it, for now, in a self contained form, postponing to the next section a more detailed discussion on the topic.
@@
We address the existence and uniqueness of this case of the asymmetric LYZ body in Section 2.1.
@@
Concerning the asymptotics for $p\to\infty$, we will see that $R_{\infty}^{m}f$ may be a compact, convex set, but it may also be $\mathbb{R}^{nm}$. Additionally, the limit $\lim_{p\to\infty}\Gamma\left(1+p\right)^{-\frac{1}{p}}R^{m}_{p}f$ will always be a compact, convex set, but it may be a singleton containing the origin, as the following statement shows. Recall the digamma function $\psi(t)=\Gamma^{\prime}(t)/\Gamma(t)$.
@@
Fix $m,n\in\mathbb{N}$. Let $K\subset\mathbb{R}^{n}$ be a convex body and set $\gamma_{K}(x)=e^{-\frac{\|x\|_{K}^{2}}{2}}$. Then, $R_{0}^{m}\gamma_{K}=\sqrt{2}\exp\left(\frac{1}{2}\psi\left(1+\frac{n}{2}\right)\right)R_{0}^{m}K$ and, for $p>0$,
$R^{m}_{p}\gamma_{K}=\sqrt{2}\left(\frac{\Gamma\left(1+\frac{n+p}{2}\right)}{\Gamma\left(1+\frac{n}{2}\right)}\right)^{\frac{1}{p}}R_{p}^{m}K.$
@@
Fix $n,m\in\mathbb{N}$. Let $f\in\operatorname{LC}_{n}$. Then, for $1<p<q<\infty$,
@@ 1.4. Functional, higher-order Rogers-Shephard inequality
Fix $n,m\in\mathbb{N}$. Let $f\in\operatorname{LC}_{n}$. Then, $\mathop{\mathrm{supp}}\nolimits(g_{f,m})\subseteq\mathbb{R}^{nm}$ contains the origin in its interior and
@@ 2.1. The LYZ body of log-concave functions
In usual applications in the literature (compare [50, 68, 49, 18]), the LYZ body of $f$ is the unique origin symmetric convex body whose surface area measure equals the even part of $\mu_{f}$. The symmetric LYZ body is a fundamental tool in establishing affine-Sobolev-type inequalities [73, 52, 36, 68, 51, 34]. Similar applications of the asymmetric body $\langle f\rangle$ can be found in [32, 33]. We emphasize that, when dealing with $f\in\operatorname{LC}_{n}$, the body $\langle f\rangle$ is not necessarily symmetric, since the level sets of $f$ are always convex and bounded and, therefore, the measure $\mu_{f}$ we retrieve in Proposition 2.3 is centered by construction, without requiring any further step.
@@ 3. Sets associated with functions
This section is dedicated to obtaining (1.25), (1.27), and Theorem 1.6 as special cases of a more general statement. We begin by recalling that a function $f:\mathbb{R}^{n}\longrightarrow\mathbb{R}_{+}$ is $s$-concave, $s>0$, if it is upper semi-continuous, not identically zero, and, for $\lambda\in[0,1]$ and $x,y\in\mathbb{R}^{n}$ such that $f(x)f(y)>0$, one has
@@
(3.13) $\begin{split}\rho_{K_{p}(g)}(\theta)&=\left(p\mathcal{M}_{\widetilde{g(\cdot\theta)}}(p)\right)^{\frac{1}{p}}\\ &=\begin{cases}\left(\frac{p}{\|g\|_{\infty}}\int_{0}^{\infty}g(r\theta)r^{p-1}\,\mathrm{d}r\right)^{\frac{1}{p}},&p>0,\\ \exp\left(\frac{1}{\|g\|_{\infty}}\int_{0}^{\infty}(-\frac{\partial}{\partial r}g(r\theta))\log(r)\;\mathrm{d}r\right),&p=0,\\ \left(\frac{p}{\|g\|_{\infty}}\int_{0}^{\infty}r^{p-1}(g(r\theta)-g(o))\;\mathrm{d}r\right)^{\frac{1}{p}},&p\in(-1,0).\end{cases}\end{split}$
@@
When $p\in(-1,0)$, an analogous requirement to (3.14) is that $g$ has maximum at the origin and the function $x\mapsto g(o)-g\left(\frac{x}{|x|^{2}}\right)$ has finite (p)(-p)-moment. Indeed, from a change of variables
$p\int_{0}^{\infty}r^{p-1}(g(r\theta)-g(o))\mathrm{d}\,r=|p|\int_{0}^{\infty}r^{|p|-1}\left(g(o)-g\left(\frac{\theta}{r}\right)\right)\mathrm{d}\,r,$
@@
$K_{\infty}(g):=\lim_{p\to\infty}K_{p}(g)=\mathop{\mathrm{supp}}\nolimits(g),$
@@ 4.2. The Functional Covariogram
Fix $m,n\in\mathbb{N}$. Let $f\in\operatorname{LC}_{n}$. Then, for every $\bar{\theta}\in\mathbb{S}^{nm-1}$,
@@
Fix $n,m\in\mathbb{N}$. Let $f\in\operatorname{LC}_{n}$. Then, $g_{f,m}$ is log-concave and, for a fixed $\bar{x}\in\mathbb{R}^{nm}$,
@@ 4.3. The approach by radial mean bodies
Clearly, the definition is well-posed, since the restriction of a summable log-concave function on $\mathbb{R}^{n}$ to a line is a summable one-dimensional log-concave function. Define the function $f_{y,m}:\mathbb{R}^{nm}\longrightarrow\mathbb{R}_{+}$ by
$f_{y,m}(\bar{x})=\min_{0\leq i\leq m}\left\{f(y-x_{i})\right\},$
@@
We finish this section by characterizing the radial mean bodies of a particular class of functions which contains those of Corollary 1.5. Recall that $R^{m}_{p}K=K_{p}(g_{K,m})$, i.e. its radial function is similar to (4.7) and (4.8), but with $\|f\|_{1}$ and $g_{f,m}$ replaced with $\mathrm{vol}_{n}(K)$ and $g_{K,m}$ respectively.
@@
Fix $m,n\in\mathbb{N}$. Let $\varphi:\mathbb{R}_{+}\longrightarrow\mathbb{R}_{+}$ be a measurable function with finite (n1)(n-1)th moment. Suppose further that $\varphi$ has maximum at the origin and decreases monotonically to zero at $\infty$. Fix $K\subset\mathbb{R}^{n}$ a convex body and define the function $f:\mathbb{R}^{n}\longrightarrow\mathbb{R}_{+}$ by $f=\varphi(\|\cdot\|_{K})$.
Then, if $p>0$ and $\varphi$ also has finite (n+p1)(n+p-1)th moment, one has
$R^{m}_{p}f=\left(\frac{n+p}{n}\frac{\int_{0}^{\infty}\varphi(t)t^{n+p-1}\mathrm{d}t}{\int_{0}^{\infty}\varphi(t)t^{n-1}\;\mathrm{d}t}\right)^{\frac{1}{p}}R_{p}^{m}K,$
while, if $p=0$ and $\varphi(t)t^{n-1}\log(t)$ is integrable on $\mathbb{R}_{+}$, one obtains
$R^{m}_{0}f=e^{\frac{1}{n}}\exp\left(\frac{\int_{0}^{\infty}\varphi(t)t^{n-1}\log(t)\;\mathrm{d}t}{\int_{0}^{\infty}\varphi(t)t^{n-1}\;\mathrm{d}t}\right)R_{0}^{m}K.$
@@
First, we compute $g_{f,m}$. Since $\varphi$ is monotone decreasing, it is differentiable almost everywhere. Therefore, for every $\bar{x}\in\mathbb{R}^{nm},$ it holds
$g_{f,m}(\bar{x})$ $=\int_{\mathbb{R}^{n}}\min_{0\leq i\leq m}\varphi(\|y-x_{i}\|_{K})\;\mathrm{d}y=\int_{\mathbb{R}^{n}}\varphi\left(\max_{0\leq i\leq m}\|y-x_{i}\|_{K}\right)\;\mathrm{d}y$
$=\int_{\mathbb{R}^{n}}\int_{\max_{0\leq i\leq m}\|y-x_{i}\|_{K}}^{\infty}(-\varphi^{\prime}(t))\;\mathrm{d}t\;\mathrm{d}y$
$=\int_{0}^{\infty}(-\varphi^{\prime}(t))\mathrm{vol}_{n}\left(tK\cap\bigcap_{i=1}^{m}\left(tK+x_{i}\right)\right)\;\mathrm{d}t$
$=\int_{0}^{\infty}(-\varphi^{\prime}(t))t^{n}\mathrm{vol}_{n}\left(K\cap\bigcap_{i=1}^{m}\left(K+t^{-1}x_{i}\right)\right)\;\mathrm{d}t$
$=\int_{0}^{\infty}(-\varphi^{\prime}(t))t^{n}g_{K,m}(t^{-1}\bar{x})\;\mathrm{d}t.$
@@
(4.10) $\begin{split}\int_{\mathbb{R}^{n}}\varphi(\|x\|_{K})\;\mathrm{d}x&=\int_{\mathbb{S}^{n-1}}\int_{0}^{\infty}\varphi(r\|\theta\|_{K})r^{n-1}\,\mathrm{d}r\;\mathrm{d}\theta\\ &=\left(n\int_{0}^{\infty}\varphi(r)r^{n-1}\;\mathrm{d}r\right)\left(\frac{1}{n}\int_{\mathbb{S}^{n-1}}\|\theta\|_{K}^{-n}\;\mathrm{d}\theta\right)\\ &=\mathrm{vol}_{n}(K)\left(n\int_{0}^{\infty}\varphi(r)r^{n-1}\;\mathrm{d}r\right).\end{split}$
We compute the radial function of $R^{m}_{p}f$ when $p>0$: for every $\bar{\theta}\in\mathbb{S}^{nm-1}$,
@@
$=\frac{p}{\|f\|_{1}}\int_{0}^{\infty}\int_{0}^{\infty}(-\varphi^{\prime}(t))t^{n}g_{K,m}(rt^{-1}\bar{\theta})r^{p-1}\;\mathrm{d}t\;\mathrm{d}r$
$=\frac{p}{\|f\|_{1}}\int_{0}^{\infty}\int_{0}^{\infty}(-\varphi^{\prime}(t))t^{n+p}g_{K,m}(r\bar{\theta})r^{p-1}\;\mathrm{d}t\;\mathrm{d}r$
$=\left(\frac{p}{\mathrm{vol}_{n}(K)}\int_{0}^{\infty}g_{K,m}(r\bar{\theta})r^{p-1}\mathrm{d}r\right)\left(\frac{\mathrm{vol}_{n}(K)}{\|f\|_{1}}\int_{0}^{\infty}(-\varphi^{\prime}(t))t^{n+p}\mathrm{d}t\right)$
$=\rho_{R^{m}_{p}K}(\bar{\theta})^{p}\left(\frac{\int_{0}^{\infty}(-\varphi^{\prime}(t))t^{n+p}\mathrm{d}t}{n\int_{0}^{\infty}\varphi(t)t^{n-1}\;\mathrm{d}t}\right).$
The claim follows since, by integration by parts,
$\int_{0}^{\infty}(-\varphi^{\prime}(t))t^{n+p}\mathrm{d}t=(n+p)\int_{0}^{\infty}\varphi(t)t^{n+p-1}\;\mathrm{d}t.$
For $p=0$, we first take the partial derivative of $g_{f,m}(\bar{x})$: since $r\mapsto g_{K,m}(r\bar{\theta})$ is $(1/n)$-concave, it is bounded from above by an $(1/n)$-affine function, specifically $g_{K,m}(r\bar{\theta})\leq\mathrm{vol}_{n}(K)\left(1-\frac{r\|\bar{\theta}\|_{\Pi^{\circ,m}K}}{n\mathrm{vol}_{n}(K)}\right)_{+}^{n}$. Therefore, we may apply the Leibniz rule and obtain, for every $r>0$ and $\bar{\theta}\in\mathbb{S}^{nm-1},$
(4.11) $\frac{\partial}{\partial r}g_{f,m}(r\bar{\theta})=\int_{0}^{\infty}(-\varphi^{\prime}(t))t^{n}\left(\frac{\partial}{\partial r}g_{K,m}(t^{-1}r\bar{\theta})\right)\;\mathrm{d}t.$
Inserting our choice of $f$ into (4.8) when $p=0$, (4.11) and (4.10) yield
$\rho_{R^{m}_{p}f}(\bar{\theta})=\exp\left(\int_{0}^{\infty}\left(-\frac{\partial}{\partial r}\!\frac{g_{f,m}(r\bar{\theta})}{\|f\|_{1}}\right)\log(r)\,\mathrm{d}r\right)$
$=\exp\left(\int_{0}^{\infty}\left(\frac{(-\varphi^{\prime}(t))t^{n}}{n\int_{0}^{\infty}\varphi(t)t^{n-1}\;\mathrm{d}t}\right)\int_{0}^{\infty}\left(-\frac{\partial}{\partial r}\frac{g_{K,m}(t^{-1}r\bar{\theta})}{\mathrm{vol}_{n}(K)}\right)\log(r)\,\mathrm{d}r\mathrm{d}t\right).$
We next introduce the variable substitution $r=st$. Adjusting both the integral and derivative in $r$, we obtain, by inserting $\rho_{R_{0}^{m}K}$ and using
$\int_{0}^{\infty}\left(-\frac{\partial}{\partial s}\frac{g_{K,m}(s\bar{\theta})}{\mathrm{vol}_{n}(K)}\right)\;\mathrm{d}s=1,$
the formula
$\rho_{R^{m}_{p}f}(\bar{\theta})\!=\!\exp\left(\!\int_{0}^{\infty}\!\!\!\left(\frac{(-\varphi^{\prime}(t))t^{n}}{n\int_{0}^{\infty}\varphi(t)t^{n-1}\;\mathrm{d}t}\right)\int_{0}^{\infty}\left(-\frac{\partial}{\partial s}\frac{g_{K,m}(s\bar{\theta})}{\mathrm{vol}_{n}(K)}\right)\log(st)\,\mathrm{d}s\mathrm{d}t\right)$
$=\exp\left(\frac{\int_{0}^{\infty}(-\varphi^{\prime}(t))t^{n}\log(t)\;\mathrm{d}t}{n\int_{0}^{\infty}\varphi(t)t^{n-1}\;\mathrm{d}t}\right)\exp\left(\left(\frac{\int_{0}^{\infty}(-\varphi^{\prime}(t))t^{n}\;\mathrm{d}t}{n\int_{0}^{\infty}\varphi(t)t^{n-1}\;\mathrm{d}t}\right)\log\rho_{R_{0}^{m}K}(\bar{\theta})\right).$
Finally, by integration by parts, we obtain
$\rho_{R^{m}_{0}f}(\bar{\theta})=e^{\frac{1}{n}}\exp\left(\frac{\int_{0}^{\infty}\varphi(t)t^{n-1}\log(t)\;\mathrm{d}t}{\int_{0}^{\infty}\varphi(t)t^{n-1}\;\mathrm{d}t}\right)\rho_{R_{0}^{m}K}(\bar{\theta}).$
A few remarks are in order. Firstly, if $f=e^{-\|\cdot\|_{K}},$ i.e. $\varphi(t)=e^{-t}$, we obtain $R_{0}^{m}f=e^{\psi(n+1)}R_{0}^{m}K$, and, for $p>0$
$R^{m}_{p}f=\left(\frac{\Gamma(n+p+1)}{\Gamma(n+1)}\right)^{\frac{1}{p}}R_{p}^{m}K.$
Using $\Gamma(n+1)^{-\frac{1}{p}}\to 1$ as $p\to\infty$, we have by Stirling’s approximation
@@
$=\lim_{p\to\infty}\left(\frac{\Gamma(n+p+1)}{\Gamma(n+1)\Gamma(p+1)}\right)^{\frac{1}{p}}R_{p}^{m}K=R_{\infty}^{m}K=D^{m}(K).$
Next, Corollary 1.5 then follows from Theorem 4.7 with the choice $\varphi(t)=e^{-\frac{t^{2}}{2}}$ and an application of Stirling’s approximation.
Finally, replicating the proof of Theorem 4.7 when $p\in(-1,0)$ yields
$\rho_{R^{m}_{p}f}(\bar{\theta})$
$=\left(\!\frac{p}{\mathrm{vol}_{n}(K)}\!\int_{0}^{\infty}\!\!\!\left(\!\!\!\left(\!\frac{n+p}{n}\frac{\int_{0}^{\infty}\varphi(t)t^{n+p-1}\mathrm{d}t}{\int_{0}^{\infty}\varphi(t)t^{n-1}\;\mathrm{d}t}\right)\!g_{K,m}(r\bar{\theta})-\!\!\mathrm{vol}_{n}(K)\!\!\right)r^{p-1}\mathrm{d}r\right)^{\frac{1}{p}}.$
The most direct and seemingly only way to relate this to $R^{m}_{p}K$ is to choose $\varphi(t)=Ce^{-\frac{n}{|p|}t^{|p|}}$ for some $C>0$, in which case the ratio of terms adjacent to $g_{K,m}$ is $1$. We outline this choice as a corollary of Theorem 4.7. We use that $C$ is a free parameter and set $C=e^{\frac{n}{|p|}}$, creating $\varphi(t)=e^{-\frac{n}{|p|}\left(t^{|p|}-1\right)}$. As $p\to 0$, this function converges to $t^{-n}$. We see that this limiting function does not satisfy the integrability requirements of the $p=0$ case of Theorem 4.7. Therefore, in the following corollary, we define a family of functions $f_{p,K}=\varphi\left(\|\cdot\|_{K}\right)$ and then arrive at $p=0$ by taking the limit as $p\to 0^{-}$ and as $p\to 0^{+}$, deriving in this way two bodies. This is because the formula for the coefficients has a jump discontinuity at $p=0$.
Corollary 4.8.
Fix $m,n\in\mathbb{N}$. Let $K\subset\mathbb{R}^{n}$ be a convex body and set
$f_{p,K}(x)=\begin{cases}e^{-\frac{n}{|p|}\left(\|x\|_{K}^{|p|}-1\right)},&p>-1,p\neq 0,\\ \|x\|_{K}^{-n},&p=0.\end{cases}$
Then,
$R^{m}_{p}f_{p,K}=\begin{cases}R_{p}^{m}K,&p\in(-1,0),\\ R_{0}^{m}K,&p=0^{-},\\ e^{1/n}R_{0}^{m}K,&p=0^{+},\\ \left(1+\frac{p}{n}\right)^{\frac{1}{p}}R_{p}^{m}K,&p>0.\end{cases}$